OpenStudy (anonymous):

If x > 0 , then 2 percent of 5 percent of 3x equals what percent of x ?

OpenStudy (anonymous):

0.3 % of X

OpenStudy (anonymous):

5 percent of 3x -> (5/100) * 3x = 15x/100 = 0.15x 2 percent of this -> (2/100)*0.15x = (0.3x/100) = 0.003x ((percent of x) * x) / 100 = 0.003x percent of x = 0.3